Country A and country B both have the production function Y = F(K, L) = K1/2L1/2.

Does this production function have constant returns to scale? Explain.

What is the per-worker production function, y = f(k)? [Hint: use the solow model]  


1
Expert's answer
2021-11-22T09:56:56-0500

Solution:

Yes, this production function exhibits constant returns to scale.

Constant returns to scale require that zY = F(zK, zL)

Here, we see that F(zK, zL) = (zK)1/2 (zL)1/2  = z1/2(K1/2L1/2) = z(K1/2 L1/2 )

Thus, it DOES exhibit the property of constant returns to scale.

 

The per-worker production function calculation:

Y = F (K, L) = K1/2L1/2

The per-worker production function is solved by dividing both sides of the above equation by L:

"\frac{K}{L} = (\frac{K}{L})^{1/2}"


Which you can rewrite as y = (k)1/2

The per-worker production function: y = (k)1/2